# Critical LAN and Score Tests for Mixed Fractional Models under High-Frequency Observation at $H=3/4$

### 24 Mar 2026 v1

@@ Abstract
We study the critical boundary $H=3/4$ for two mixed fractional models under high-frequency observation, namely mixed fractional Brownian motion and mixed fractional Ornstein–Uhlenbeck. For different reasons, both the supercritical normalization for $H>3/4$ and the subcritical normalization for $H<3/4$ fail at this value. We identify the exact critical first-order scales, prove the corresponding score central limit theorems, and show that once the explicit linear term in the $H$-score is removed, the resulting $(\sigma,H)$-block is already non-degenerate. Thus, in contrast with the regime $H>3/4$, the critical point is resolved by a single triangular local reparametrization. Building on this reduction, we derive the critical second-order likelihood expansion and establish LAN for both models with fully explicit leading information constants.
Motivated by a testing viewpoint from mathematical finance, we additionally derive score-type tests at the critical boundary. In particular, we formulate one-sided procedures for detecting whether the system enters the supercritical side $H>3/4$. The proposed tests are calibrated under the critical constraint $H=3/4$, use a right-tail rejection region dictated by the LAN drift, and admit fully explicit normalizations at the leading order.
@@ 1 Introduction
Mixed fractional models play an important role in financial econometrics and the statistics of stochastic processes. By superimposing a standard Brownian motion with a fractional Brownian motion, these models capture long-range dependence while preserving a semimartingale structure, thereby ruling out arbitrage opportunities in frictionless markets [4]. In recent years, high-frequency statistical inference for the mixed fractional Brownian motion (mfBm) and the mixed fractional Ornstein–Uhlenbeck (mfOU) process has attracted considerable attention.
For fractional noise models, the low frequency behaviour determines the asymptotic properties of estimators. [1] established local asymptotic normality (LAN) for fractional Gaussian noise under high-frequency observations, revealing the role of the Hurst exponent $H$. [2] extended the LAN theory to mixed fractional Brownian motion and showed that for $H>3/4$ two non-diagonal transformations are required to regularise the parameter space, whereas for $H<3/4$ a single transformation suffices. [3] obtained analogous results for the mixed fractional OU process and further demonstrated that the $\alpha$-direction is asymptotically orthogonal to the $(\sigma,H)$-block under the critical scaling.
Nevertheless, all the above works exclude the boundary $H=3/4$. At this critical point, the low frequency singularity of the spectral density achieves a precise balance between signal and noise, causing both the supercritical and subcritical normalisations to break down. Specifically, when $H=3/4$ the transformed $(\sigma,H)$-block is no longer singular, but the leading traces acquire an extra logarithmic factor, necessitating a novel scaling. The statistical inference at this critical boundary has remained unexplored in the existing literature.
From a practical perspective, detecting whether the system enters the supercritical region $H>3/4$ is of particular interest in continuous time finance, as it signals stronger long-range dependence and may affect risk management and derivative pricing. Consequently, developing a testing procedure that is calibrated at the boundary $H=3/4$ and has a clear directional interpretation becomes a natural objective.
This paper fills the above gaps. We identify the exact first-order scales at $H=3/4$, prove central limit theorems for the transformed score vectors, and establish a complete LAN expansion whose information matrix is given by fully explicit constants for two models
  1. (i)
    mixed fractional Brownian motion (mfBm),
  2. (ii)
    mixed fractional Ornstein–Uhlenbeck (mfOU).
In many continuous time financial models, semimartingale type restrictions are invoked to rule out arbitrage under frictionless trading. Mixed fractional models exhibit a pronounced change of local asymptotic structure around the boundary $H=3/4$. Thus this boundary is a natural point at which to formulate statistical tests for regime identification.
In this paper, we focus on the one-sided testing problem
$H_{0}:\ H\leq\frac{3}{4}\qquad\text{vs}\qquad H_{1}:\ H>\frac{3}{4},$
which is natural from the supercritical viewpoint. Since the score statistics are constructed at $H=3/4$, the sign of the local LAN drift in the $H$-direction leads to a right-tailed rejection region. For completeness, we also record the corresponding two-sided boundary test.
Notation.
  • For an integrable symbol $g$ on $\pi,\pi]$, let
    $\mathcal{T}_{n}(g):=\bigl(\widehat{g}(i-j)\bigr)_{1\leq i,j\leq n},\qquad\widehat{g}(k):=\frac{1}{2\pi}\int_{-\pi}^{\pi}g(\lambda)e^{-ik\lambda}\,\mathrm{d}\lambda.$
    We reserve $T_{n}=n\Delta_{n}$ for the observation horizon; this removes the ambiguity between the time horizon and the Toeplitz operator. If $A_{n}$ is a real symmetric matrix and $Z_{n}\sim\mathcal{N}(0,I_{n})$, write
    $Q_{n}(A_{n}):=Z_{n}^{\top}A_{n}Z_{n}-\mathrm{tr}(A_{n}).$ (1)
  • Throughout, we assume that
    $\Delta_{n}\downarrow 0,\qquad T_{n}:=n\Delta_{n}\to\infty,\qquad L_{n}:=\log(1/\Delta_{n}).$
    All matrices and constants are evaluated at the critical point $H=3/4$.
@@
The paper is organized as follows. Section 2 introduces the mfBm and mfOU models, derives the critical low frequency constants and covariance matrices at $H=3/4$, and establishes the exact score decompositions and critical score profiles used throughout this paper. Section 3 states the critical score CLTs and the LAN theorems for the mfBm and mfOU models. Sections 4 and 5 prove these results. Section 6 develops score-type tests at the critical boundary $H=3/4$. Section 7 reports a simulation study for the feasible mfBm test. Appendix A collects the low frequency reductions, the fully explicit constants and the higher-order remainder bounds.
2 Preliminaries
In this section, we collect the model setup, notation and critical quantities needed for the analysis at the critical point $H=3/4$. We first introduce the mfBm and mfOU models, together with their low frequency spectral behavior at the critical point. We then define the corresponding critical covariance constants and matrices, derive the exact Gaussian score decompositions, and define the critical score profiles.
@@ 2.1 The mfBm model
$K_{\mathrm{m}}=c_{3/4},\qquad\beta_{\mathrm{m}}=2\psi\!\left(\frac{5}{2}\right)-\pi,$
@@ 2.2 The mfOU model
Let $f_{\Delta}^{\mathrm{o}}(\lambda;\sigma,H,\alpha)$ be the sampled spectral density of the stationary mfOU model (see [3]). At the continuous time frequencies, the stationary mfOU solution has a smooth OU transfer factor multiplying the same fractional singular coefficient $c_{H}=\Gamma(2H+1)\sin(\pi H)$ as in (2). Consequently, the sampled spectral density inherits the same low frequency singular coefficient and logarithmic derivative in the $H$-direction. More precisely, under the normalization used throughout the paper,
@@ 2.3 Critical covariance constants
Define
@@
Hence the critical $(\sigma,H)$-block is already non-degenerate after the first transformation. The low frequency reductions and the explicit one-dimensional integrals leading to (8)–(9) are collected in Appendix A.
@@ 2.4 Exact score functions
where $Q_{n}(\cdot)$ is defined in equation (1).
@@
We now rewrite the raw score functions at the critical boundary $H=3/4$ into exact decompositions that isolate the singular $\sigma$-direction in the $H$-score. This yields the transformed $H$-scores. For mfBm define
@@
The positive sign in the definition of $D_{n}^{\mathrm{m}}$ comes from replacing $\log\Delta_{n}$ by $L_{n}$; this sign convention keeps the critical decomposition unambiguous.
@@
The additional factor $2L_{n}$ in (20) survives after removing the explicit $\sigma L_{n}\,S_{\sigma,n}$ term; it comes from the derivative of the low frequency singularity itself.
@@
3 Main Results
We now state the central limit theorems for the transformed score vectors in equations (15) and (16) at the critical boundary $H=3/4$. Throughout this subsection, let
@@ 3 Main Results
and all score matrices and score functions are evaluated at the critical parameter point. For the mfBm model we write
@@
and for the mfOU model,
@@
Recall that the $H$-score admits the exact decomposition
$S^{m}_{H,n}=-\sigma L_{n}S^{m}_{\sigma,n}+R^{m}_{H,n},\qquad S^{o}_{H,n}=-\sigma L_{n}S^{o}_{\sigma,n}+R^{o}_{H,n},$
where $R^{m}_{H,n}$ and $R^{o}_{H,n}$ denote the transformed $H$-scores after removing the explicit singular linear term in the $\sigma$-direction.
The next two theorems show that, after normalization by the exact critical rates, the transformed score vectors are asymptotically Gaussian.
@@
where $\Gamma_{\mathrm{o}}^{\mathrm{crit}}$ is defined by equation (11). Hence, at the critical point, the $(\sigma,H)$-block and the $\alpha$-direction are asymptotically orthogonal at the leading order.
The previous theorems are formulated in terms of the transformed $H$-score $R_{H,n}$ rather than the raw $H$-score $S_{H,n}$. By (15) and (16), the two score functions are related by the following triangular changes of coordinates.
@@
These matrices encode exactly the critical local scales of the parameters and the first-order coupling between the $\sigma$- and $H$-directions.
@@
4 Proof of the Score CLTs
This section proves the score central limit theorems. We first establish the critical trace asymptotics for the score matrices and then combine them with the Gaussian quadratic-form CLT (see lemma 13) to derive the limiting normal laws.
@@ 4.1 Critical trace asymptotics
$\left\lVert C_{n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}=O(1),\qquad\left\lVert D_{n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}=O(L_{n}).$ (25)
@@
are reduced to the corresponding one-dimensional critical integrals via the standard Toeplitz trace reduction (see [2]).
@@
For $C_{n}^{m}$, the operator bound follows from the boundedness of the corresponding reduced symbol. For $D_{n}^{m}$, however, the reduced profile $g_{H,n}$ is not bounded near the origin because of the logarithmic term. Hence the bound $D_{n}^{m}||_{op}=O(L_{n})$ requires a separate argument at the level of the original sandwiched symbol, which we omit here and prove below. ∎
@@
$\left\lVert C_{n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}=O(1),\qquad\left\lVert D_{n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}=O(L_{n}),\qquad\left\lVert A_{n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}=O(1).$ (30)
@@
For the $(\sigma,H)$-block, the critical rescaling of the sampled mfOU spectral density yields the same integrals as in the mfBm case (see [3]), with $\eta_{\mathrm{m}}$ replaced by $\eta_{\mathrm{o}}$ (see Appendix A for the explicit low frequency reduction). This proves (26)–(28).

### 02 Apr 2026 v2

# Boundary Inference for Mixed Fractional Models under High-Frequency Observation Critical LAN and Score Tests at $H=3/4$

@@ Abstract
We study boundary inference at $H=3/4$ for mixed fractional Brownian motion and mixed fractional Ornstein–Uhlenbeck models under high-frequency observation. This boundary is economically important because it separates the critical and supercritical regimes of mixed fractional dynamics.
We make three contributions. First, we identify the exact critical first-order scaling and show that, after removing the explicit linear component in the $H$-score, the transformed $(\sigma,H)$ block is already non-degenerate. Second, we establish critical score central limit theorems (CLT) and derive local asymptotic normality (LAN) with fully explicit leading information constants for both models. Third, we construct boundary-calibrated one-sided score tests for detecting entry into the supercritical region $H>3/4$ and discuss feasible implementation through restricted nuisance estimation. Monte Carlo evidence shows that the feasible statistic has the correct directional power but conservative null calibration. Finally, an intraday illustration on one-minute SPY data finds no persistent evidence in favor of $H>3/4$.
Keywords: mixed fractional Brownian motion; mixed fractional Ornstein–Uhlenbeck process; boundary inference; local asymptotic normality; score test; high-frequency data.
JEL classification: C12; C22; C58.
@@ 1 Introduction
Mixed fractional models sit naturally at the intersection of continuous-time probability and financial econometrics. On the one hand, they enrich Brownian dynamics by adding a fractional component capable of carrying persistent dependence; on the other hand, they remain close enough to Gaussian likelihood theory to support high-frequency inference. In empirical work, such models are appealing whenever one wants to describe serially dependent microstructure distortions, persistent volatility signals, or reduced-form deviations from the benchmark semimartingale paradigm. For mixed fractional Brownian motion, however, the semimartingale interpretation is known to hinge on the critical threshold $H=3/4$ in Cheridito (2001). The point $H=3/4$ is therefore not merely a mathematical edge case. It is the place where the probabilistic status of the model, and hence its econometric interpretation, changes in a decisive way.
That question has become more rather than less relevant in recent years. The modern rough-volatility literature has made fractional behavior a central empirical theme in finance, with influential contributions in Quantitative Finance and related outlets emphasizing rough volatility surfaces, short-maturity asymptotics, and the high-frequency evidence for very small Hurst exponents, see, e.g., Bayer et al. (2016), Fukasawa (2017), Gatheral et al. (2018), and Fukasawa (2021). At the same time, the Journal of Econometrics has developed a complementary line of work that treats fractional and rough dynamics as econometric objects to be estimated, tested, and used in forecasting. Examples include the fOU realized-volatility framework of Wang et al. (2023), the GMM-based roughness estimation strategy of Bolko et al. (2023), and the approximate Whittle likelihood analysis of Shi et al. (2024). Much of this literature focuses, understandably, on rough regimes with $H<1/2$ or on estimation away from critical boundaries. Yet if mixed fractional models are to be used as diagnostic or structural tools, then the opposite side of the problem matters as well: one needs a statistically disciplined way to decide whether the data are still compatible with the critical or subcritical side of the model, or whether they have crossed into the supercritical region where the semimartingale logic becomes substantially more delicate.
This is precisely where the present paper enters. Existing LAN results for high-frequency fractional models do not resolve the boundary point $H=3/4$. For fractional Gaussian noise, Brouste and Fukasawa (2018) established LAN under high-frequency observation. For mixed fractional Brownian motion and mixed fractional Ornstein–Uhlenbeck processes, Cai and Shang (2026) and Cai et al. (2026) showed that the local geometry differs sharply across the subcritical and supercritical regions. But the critical case itself is excluded from those analyses. At $H=3/4$, both off-boundary normalizations fail. The leading traces acquire an additional logarithmic factor, the raw $H$-score develops an explicit singular coupling with the nuisance direction, and neither side of the existing theory can simply be pasted onto the boundary.
Our main observation is that the critical point is singular only before the right coordinates are chosen. Once the explicit linear term in the $H$-score is removed, the remaining $(\sigma,H)$ block is already non-degenerate at $H=3/4$. This fact turns out to be enough to recover an exact critical normalization, to write down the limiting covariance structure with explicit constants, and to obtain a single triangular local reparametrization for both mixed fractional experiments. In that sense, the paper does not merely fill in a missing boundary case. It shows that the boundary has its own coherent local asymptotic geometry, and that this geometry leads directly to a usable score-based inferential problem.
The testing question that emerges is economically and econometrically transparent. We study
$H_{0}:\ H\leq\frac{3}{4}\qquad\text{versus}\qquad H_{1}:\ H>\frac{3}{4},$
so the null collects the critical and subcritical configurations while the alternative corresponds to entry into the supercritical regime. When mixed fractional models are interpreted as reduced-form descriptions of persistent high-frequency dependence, rejection in favor of $H>3/4$ can be read as evidence of an exceptionally persistent regime, one that is harder to reconcile with the semimartingale benchmark underlying much of continuous-time financial econometrics. Conversely, non-rejection supports staying on the critical side of the boundary, where the local asymptotic structure is fundamentally different. This is the sense in which our score tests should be understood: they are not only formal consequences of LAN expansions, but boundary diagnostics for a class of models that sits directly beside the rough-volatility and high-frequency econometrics literatures.
The contribution of the paper is therefore theoretical in method but econometric in orientation. We identify the exact critical first-order scales, prove Gaussian limits for transformed score vectors in both the mfBm and mfOU experiments, derive explicit LAN expansions with fully specified leading constants, and use the resulting central sequences to construct one-sided score tests calibrated at the boundary. We also discuss feasible implementation through nuisance estimation under the restriction $H=3/4$, report an expanded Monte Carlo analysis of the feasible mfBm statistic, and close with a compact intraday empirical illustration built around that same boundary score.
@@
The rest of the paper is organized as follows. Section 2 introduces the mfBm and mfOU models, derives the critical low-frequency constants and covariance matrices at $H=3/4$, and establishes the exact score decompositions that drive the subsequent analysis. Section 3 states the transformed score CLTs, the critical LAN expansions. Sections 4 and 5 contain the technical proofs. Section 6 develops the boundary score tests and explains feasible implementation under nuisance estimation. Section 7 reports the expanded Monte Carlo evidence. Section 8 turns to empirical implementation of the exact feasible score on genuinely intraday data. The appendices collect the low-frequency reductions, explicit constants and higher-order remainder bounds.
2 Model setup, critical objects and score decomposition
We now introduce the objects that will carry the boundary analysis. Rather than introducing the models and the asymptotic ingredients in isolation, it is helpful to keep the inferential goal in view from the outset. Everything that follows is organized around one question: what remains of the Gaussian score structure exactly at $H=3/4$, once the usual off-boundary normalizations have ceased to be informative? For that reason, the section moves from the model definitions to the low-frequency behavior that drives the critical rates, and then from those rates to the score decomposition that eventually makes the boundary experiment tractable.
@@ 2.1 The mfBm model
$K_{\mathrm{m}}:=c_{3/4},\qquad\beta_{\mathrm{m}}:=2\psi\!\left(\frac{5}{2}\right)-\pi,$
@@
Proposition 1 identifies the low frequency behavior of the fGn spectral density and its derivative in the Hurst direction. In particular, the constants $K_{m}$ and $\beta_{m}$ extracted here will enter the critical score profiles, the covariance matrix, and the subsequent LAN analysis.
@@ 2.2 The mfOU model
Let $f_{\Delta}^{\mathrm{o}}(\lambda;\sigma,H,\alpha)$ be the sampled spectral density of the stationary mfOU model (see Cai et al. (2026)). At the continuous time frequencies, the stationary mfOU solution has a smooth OU transfer factor multiplying the same fractional singular coefficient $c_{H}=\Gamma(2H+1)\sin(\pi H)$ as in (1). Consequently, the sampled spectral density inherits the same low frequency singular coefficient and logarithmic derivative in the $H$-direction. More precisely, under the normalization used throughout the paper,
@@ 2.3 Critical covariance constants
This subsection collects the critical covariance constants that determine the covariance matrices in the score CLTs at $H=3/4$. These quantities also enter the information matrices appearing in the subsequent LAN expansions and score tests. Define
@@
Hence the critical $(\sigma,H)$-block is already non-degenerate after the first transformation. Its connection with the transformed score will be made explicit in the next subsection. The low frequency reductions and the explicit one-dimensional integrals leading to (7)–(8) are collected in Appendix A.
@@ 2.4 Exact score functions
where $Q_{n}(A):=Z_{n}^{\top}AZ_{n}-\mathrm{tr}(A)$ denotes the centered Gaussian quadratic form.
@@
We now rewrite the raw score functions at the critical boundary $H=3/4$ into exact decompositions that isolate the singular $\sigma$-direction in the $H$-score. We refer to this step as the first transformation. This yields the transformed $H$-scores. For mfBm define
@@
The positive sign in the definition of $D_{n}^{\mathrm{m}}$ comes from replacing $\log\Delta_{n}$ by $L_{n}$, this sign convention keeps the critical decomposition unambiguous.
Equation (14) also defines what we call the first transformation at the critical boundary. We replace the raw pair $\bigl(S_{\sigma,n}^{\mathrm{m}},S_{H,n}^{\mathrm{m}}\bigr)$ by the transformed pair $\bigl(S_{\sigma,n}^{\mathrm{m}},R_{H,n}^{\mathrm{m}}\bigr)$, thereby removing the explicit singular coupling term $\sigma L_{n}\,S_{\sigma,n}^{\mathrm{m}}$ from the $H$-score.
@@
The additional factor $2L_{n}$ in (19) survives after removing the explicit $\sigma L_{n}\,S_{\sigma,n}$ term. It comes from the derivative of the low frequency singularity itself.
@@
In short, the critical profiles $g_{\sigma}^{\bullet}(u)$, $g_{H,n}^{\bullet}(u)$ (with $\bullet\in\{\mathrm{m},\mathrm{o}\}$) and $g_{\alpha}^{\mathrm{o}}(u)$ capture the low-frequency behavior of the score matrices at the boundary $H=3/4$. They translate matrix traces into explicit integrals over the scaled frequency $u$. Consequently, all leading constants in the score CLTs and LAN information matrices are expressed in closed form via integrals of these profiles, providing a unified description of the boundary experiment.
3 Main asymptotic results
3.1 Critical score central limit theorems
With the score decomposition in hand, we can now pass from exact identities to their asymptotic meaning at the boundary $H=3/4$. Throughout this section,
@@ 3.1 Critical score central limit theorems
and all score matrices and score functions are evaluated at the critical parameter point. For the mfBm experiment we write
@@
while for the mfOU experiment we write
@@
The decisive feature proved in Section 2 is that the raw $H$-score does not need to be discarded at the boundary; it needs to be recentered in the right direction. More precisely,
$S^{m}_{H,n}=-\sigma L_{n}S^{m}_{\sigma,n}+R^{m}_{H,n},\qquad S^{o}_{H,n}=-\sigma L_{n}S^{o}_{\sigma,n}+R^{o}_{H,n}.$
The explicit $\sigma L_{n}$ term is the only first-order obstruction. Once it is peeled away, the residual score has its own critical fluctuation scale and can be handled jointly with the nuisance directions. The limit theorems below show that this transformed system has a non-degenerate Gaussian limit for both mixed fractional experiments. In other words, the boundary is singular in the original coordinates but regular enough in the transformed ones to support a full LAN analysis and, later on, a score test.
@@
where $\Gamma_{\mathrm{o}}^{\mathrm{crit}}$ is defined by equation (10). In particular, the $(\sigma,H)$-block and the $\alpha$-direction separate at the leading order, so the drift parameter does not interfere with the first-order boundary geometry.
3.2 Critical LAN expansions
Theorems 6 and 7 are deliberately written in terms of the transformed $H$-score rather than the raw one. This is not a cosmetic choice. At the boundary, the transformed coordinate system is the one in which the limit experiment becomes visible without further ad hoc projections. The following proposition records that coordinate change in matrix form and makes clear why a single triangular transformation is enough.
@@ 3.2 Critical LAN expansions
These matrices encode exactly the critical local scales of the parameters together with the first-order coupling between the $\sigma$- and $H$-directions. In this form, the local perturbation already reflects the geometry revealed by the score decomposition, and the log-likelihood admits the standard quadratic LAN representation.
@@
4 Technical proof of the score CLTs
This section establishes the score central limit theorems announced above. The argument follows the logic already suggested by the critical decomposition: the Toeplitz structure first converts the relevant quadratic forms into explicit trace asymptotics, and those asymptotics are then paired with a Gaussian quadratic-form CLT to obtain the limiting normal laws.
@@ 4.1 Critical trace asymptotics
$\left\lVert C_{n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}=O(1),\qquad\left\lVert D_{n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}=O(\log n).$ (24)
@@
are reduced to the corresponding one-dimensional critical integrals via the standard Toeplitz trace reduction (see Cai and Shang (2026)).
@@
For $C_{n}^{m}$, the operator bound follows from the boundedness of the corresponding reduced symbol. For $D_{n}^{m}$, however, the reduced profile $g_{H,n}$ is not bounded near the origin because of the logarithmic term. Hence the bound $D_{n}^{m}||_{op}=O(\log n)$ requires a separate argument at the level of the original sandwiched symbol, which we omit here and the proof follows the similar argument as in Cai and Shang (2026); Cai et al. (2026). ∎
@@
$\left\lVert C_{n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}=O(1),\qquad\left\lVert D_{n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}=O(\log n),\qquad\left\lVert A_{n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}=O(1).$ (29)
@@
For the $(\sigma,H)$-block, the critical rescaling of the sampled mfOU spectral density yields the same integrals as in the mfBm case (see Cai et al. (2026)), with $\eta_{\mathrm{m}}$ replaced by $\eta_{\mathrm{o}}$ (see Appendix A for the explicit low frequency reduction). This proves (25)–(27).